### v2

# A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials

A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials
Zhiwei Zhang
Biostatistics Innovation Group, Gilead Sciences, Foster City, California, USA
zhiwei.zhang6@gilead.com
Summary
The statistical efficiency of randomized clinical trials can be improved by incorporating information from baseline covariates (i.e., pre-treatment patient characteristics). This can be done in the design stage using stratified (permutated block) randomization or in the analysis stage through covariate adjustment. This article makes a connection between covariate adjustment and stratified randomization in a general framework where all regular, asymptotically linear estimators are identified as augmented estimators. From a geometric perspective, covariate adjustment can be viewed as an attempt to approximate the optimal augmentation function, and stratified randomization improves a given approximation by moving it closer to the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to attaching an optimal augmentation term based on the stratification factor. Under stratified randomization, adjusting for the stratification factor only in data analysis is not expected to improve efficiency, and the key to efficient estimation is incorporating new prognostic information from other covariates. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. These observations are confirmed in a simulation study and illustrated using real clinical trial data.
Key words: analysis of covariance; augmented estimator; covariate-adaptive randomization; simple randomization; standardized logistic regression; super learner
1 Introduction
Randomized clinical trials are widely considered the gold standard for evaluating the safety and effectiveness of medical treatments. It is well recognized that baseline covariates (i.e., pre-treatment patient characteristics) can be used to improve the statistical efficiency of treatment effect estimation in the analysis of trial data. This usage of baseline covariates, commonly known as covariate adjustment, has received a great deal of attention from practitioners and regulators 1. Another common usage of baseline covariates in clinical trials is stratified (permuted block) randomization 2, which aims to improve the balance of selected baseline covariates between treatment groups. In this work, we study the relationship between covariate adjustment and stratified randomization with a focus on statistical efficiency.
There is a large and growing literature on statistical methods for covariate adjustment in clinical trials 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. Under simple randomization, the semiparametric theory 13 can be used to characterize all regular, asymptotically linear estimators of common treatment effect measures such as the mean (risk) difference, the mean (risk) ratio, and the odds ratio. This characterization has motivated an augmentation approach in which an unadjusted estimator or estimating function, which does not involve covariate data, is augmented with a term based on treatment and an arbitrary function of covariates 3, 4. The optimal augmentation, which minimizes the asymptotic variance of an augmented estimator, has also been characterized. To take advantage of this optimality result, one may use a working regression model or a machine learning algorithm to estimate the optimal augmentation and use the estimated augmentation to construct an augmented estimator. The methods and results referenced in this paragraph generally assume simple randomization; they are not immediately applicable to stratified randomization.
Stratified randomization is a special case of covariate-adaptive randomization (CAR) 14, 15. Under CAR, treatment assignments for different subjects may depend on each other, and this dependence needs to be taken into account when analyzing treatment effect estimators. In recent years, considerable progress has been made in understanding the asymptotic properties of treatment effect estimators and test statistics under CAR 16, 17, 18, 19, 20, 21, 22. While some of these references pertain to specific estimators, the work of Wang et al. 21 covers a large class of M-estimators (i.e., estimators that can be regarded as solutions to estimating equations), and the work of Bannick et al. 22 covers the entire class of augmented estimators for common treatment effect measures. Loosely speaking, their results indicate that a treatment effect estimator that is asymptotically linear under simple randomization typically remains so under CAR, and its asymptotic variance may become smaller. Among the various forms of CAR studied, stratified randomization and the biased coin design 23 achieve the largest variance reduction because of their “strong balance” property 17, 19, 22. In light of its statistical efficiency, simplicity and widespread use in practice, we will focus on stratified randomization as a representative of CAR.
In this article, we make a general connection between covariate adjustment and stratified randomization and discuss its practical implications. Under the aforementioned augmentation approach, which covers virtually all estimators of practical interest, we note that covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function in a suitable function space. From this geometric perspective, stratified randomization can be interpreted as a way to improve a given approximation by projecting it into a certain subspace containing the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to adding an optimal augmentation term based on the stratification factor. This connection helps characterize efficient estimators under stratified randomization and has important practical implications:
  • Under stratified randomization, adjusting for the stratification factor only in the analysis of trial data is not expected to improve statistical efficiency. Adjusting for the stratification factor is common practice in clinical trials and consistent with regulatory guidelines 24, 25, 26, 1. However, the key to efficient estimation is incorporating new prognostic information from other baseline covariates than the stratification factor (whose prognostic information has already been utilized in stratified randomization). Such covariates may include continuous covariates underlying the stratification factor as well as additional covariates unrelated to the stratification factor.
  • In designing a trial with stratified randomization, there is often a conflict between the desire to include all important covariates in the stratification and the need to avoid having too many strata, some of which may be quite small. In light of the connection between stratified randomization and covariate adjustment, it is not essential to include all important covariates in the stratification, as covariate adjustment provides another opportunity to make use of their prognostic information. This flexibility can be helpful in practice, especially for covariates that are time-consuming to measure.
To complement the above considerations based on asymptotic theory, we report a simulation study comparing different estimation methods under simple versus stratified randomization. Our simulation results confirm that adjusting for the stratification factor alone does not improve efficiency under stratified randomization, that large efficiency gains can be obtained by incorporating other covariates than the stratification factor, and that stratified randomization makes little difference to estimators that adjust for all covariates. Among the different methods for covariate adjustment, a nonparametric method based on the super learner 10 performs similarly to or better than standard methods based on parametric models, depending on how well the parametric models approximate the truth.
The rest of the article is organized as follows. In Section 2, we formulate the estimation problem and describe treatment effect estimators including augmented estimators. In Section 3, we summarize asymptotic results for stratified randomization, make a connection with covariate adjustment, and discuss its practical implications. A simulation study is reported in Section 4, followed by a real example in Section 5. The article concludes with a discussion in Section 6. Technical proofs are provided in Appendix.
2 Covariate Adjustment
2.1 Estimation Problem
We start by formulating the estimation problem using potential outcomes 27. For a generic patient in the target population, let $Y(a)$ denote the potential outcome for treatment $a$, where $a=1$ for an experimental treatment and $a=0$ for a control treatment, which may be placebo or standard of care. For each $a\in\{0,1\}$, we assume $\operatorname{E}\{Y(a)^{2}\}<\infty$ and denote $\mu_{a}=\operatorname{E}\{Y(a)\}$. Suppose the treatment effect of interest is $\delta=g(\mu_{1})-g(\mu_{0})$, where $g$ is a smooth and strictly increasing function specified by the investigator. Popular choices for $g$ include identity (for continuous and other types of outcomes), logarithm (for outcomes with positive means) and logit (for binary outcomes).
Let $\boldsymbol{W}$ be a vector of available baseline covariates (i.e., pre-treatment patient characteristics) that may be associated with either or both of $Y(1)$ and $Y(0)$. Let $\boldsymbol{Z}_{i}=(Y_{i}(1),Y_{i}(0),\boldsymbol{W}_{i})$, $i=1,\dots,n$, be independent copies of $\boldsymbol{Z}=(Y(1),Y(0),\boldsymbol{W})$ arising from a random sample of $n$ subjects enrolled in a clinical trial. For the $i$th subject in the trial, let $A_{i}$ denote the randomized treatment assignment (to be discussed later) and $Y_{i}$ the actual observed outcome. Under the assumption of consistency or stable unit treatment value, we have $Y_{i}=Y_{i}(A_{i})=A_{i}Y_{i}(1)+(1-A_{i})Y_{i}(0)$, $i=1,\dots,n$. The observed data will be denoted by $\boldsymbol{O}_{i}=(\boldsymbol{W}_{i},A_{i},Y_{i})$, $i=1,\dots,n$.
2.2 Estimators
While stratified randomization is of particular interest in this article, simple randomization provides a natural framework for introducing estimators. Therefore, in the rest of this section, we assume that simple randomization is used to make treatment assignments. This means that the $A_{i}$’s are independent of each other and of the $\boldsymbol{Z}_{i}$’s, with $\operatorname{P}(A_{i}=1)\equiv\pi\in(0,1)$. It follows that the $\boldsymbol{O}_{i}$’s are independent copies of $\boldsymbol{O}=(\boldsymbol{W},A,Y)$, where $A$ is Bernoulli with $\operatorname{P}(A=1|\boldsymbol{Z})=\operatorname{P}(A=1)=\pi$ and $Y=Y(A)=AY(1)+(1-A)Y(0)$.
It is straightforward to estimate $\delta$ with the empirical estimator $\widehat{\delta}_{\text{emp}}=g\left(\overline{Y}_{1}\right)-g\left(\overline{Y}_{0}\right)$, where $\overline{Y}_{a}=\{\sum_{i=1}^{n}I(A_{i}=a)\}^{-1}\sum_{i=1}^{n}I(A_{i}=a)Y_{i}$, $a=0,1$, and $I(\cdot)$ is the indicator function. It is well known that $\widehat{\delta}_{\text{emp}}$ is consistent and asymptotically normal in the sense that $\sqrt{n}(\widehat{\delta}_{\text{emp}}-\delta)$ converges to a zero-mean normal distribution. Its (scaled) asymptotic variance is given by $\operatorname{Var}\{\psi(\boldsymbol{O})\}$, where
$\psi(\boldsymbol{O})=g^{\prime}(\mu_{1})\frac{A(Y-\mu_{1})}{\pi}-g^{\prime}(\mu_{0})\frac{(1-A)(Y-\mu_{0})}{1-\pi}$
and $g^{\prime}$ is the derivative function of $g$. The empirical estimator may be inefficient as it does not make use of the available information in $\boldsymbol{W}$.
There are many covariate-adjusted estimators of $\delta$ that incorporate covariate information for improved efficiency. Here we describe one such estimator based on a standard regression model:
$\operatorname{E}(Y|A,\boldsymbol{V})=h\left((1,A,\boldsymbol{V}^{\text{T}},A\boldsymbol{V}^{\text{T}})\mbox{\boldmath${\beta}$}\right),$ (1)
where $\boldsymbol{V}$ is a vector-valued function of $\boldsymbol{W}$, $h$ a specified inverse link function, and ${\beta}$ an unknown parameter vector. The interaction term in model (1) is optional in practice but necessary in the subsequent discussion. Suppose ${\beta}$ is estimated by solving the estimating equation
$\sum_{i=1}^{n}\left\{Y_{i}-h(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})\mbox{\boldmath${\beta}$}\right\}(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})^{\text{T}}=\boldsymbol{0}.$ (2)
[…]
$\widehat{\mu}_{\text{reg},a}=\frac{1}{n}\sum_{i=1}^{n}h\left((1,a,\boldsymbol{V}_{i}^{\text{T}},a\boldsymbol{V}_{i}^{\text{T}})\widehat{}\mbox{\boldmath${\beta}$}\right),\qquad a=0,1.$
Examples of the regression estimator include the analysis of covariance (ANCOVA) estimator for continuous outcomes 3 and the standardized logistic regression estimator for binary outcomes 6. It is worth noting that, whether model (1) is correct or not, $\widehat{\delta}_{\text{reg}}$ is consistent for $\delta$ and asymptotically normal with asymptotic variance $\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{reg}}(\boldsymbol{W})\}$, where
$b_{\text{reg}}(\boldsymbol{W})=\frac{g^{\prime}(\mu_{1})\left\{h\left((1,1,\boldsymbol{V}^{\text{T}},\boldsymbol{V}^{\text{T}})\mbox{\boldmath${\beta}$}^{*}\right)-\mu_{1}\right\}}{\pi}+\frac{g^{\prime}(\mu_{0})\left\{h\left((1,0,\boldsymbol{V}^{\text{T}},0\boldsymbol{V}^{\text{T}})\mbox{\boldmath${\beta}$}^{*}\right)-\mu_{0}\right\}}{1-\pi}$
[…]
For full generality, we will consider a large class of augmented estimators of $\delta$ that encompass all estimators of practical interest. The augmentation approach is described in detail by the original authors 3, 4 and summarized by Zhang, Zhang and Liu 29 using notations similar to ours. In this paragraph, we give a high-level description of this approach, omitting technical details and nuances. An augmented estimator of $\delta$ may be obtained as $\widehat{\delta}_{\text{aug}}(b)=\widehat{\delta}_{\text{emp}}-n^{-1}\sum_{i=1}^{n}(A_{i}-\pi)b(\boldsymbol{W}_{i})$, where $b(\boldsymbol{W})$ is a real-valued function of $\boldsymbol{W}$ with finite variance. For any fixed $b$, $\widehat{\delta}_{\text{aug}}(b)$ is consistent and asymptotically normal with asymptotic variance $\sigma^{2}(b)=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}$. Furthermore, any estimator of $\delta$ with certain desirable properties must be asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ for some $b$ (and thus have the same asymptotic variance). For example, the regression estimator $\widehat{\delta}_{\text{reg}}$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b_{\text{reg}})$ with the same asymptotic variance $\sigma^{2}(b_{\text{reg}})$. For efficient estimation of $\delta$, it suffices to restrict attention to augmented estimators and search for the most efficient one with the smallest asymptotic variance. The optimal choice of $b$, which minimizes $\sigma^{2}(b)$, is given by
$b_{\text{opt}}(\boldsymbol{W})=\frac{g^{\prime}(\mu_{1})\{m_{1}(\boldsymbol{W})-\mu_{1}\}}{\pi}+\frac{g^{\prime}(\mu_{0})\{m_{0}(\boldsymbol{W})-\mu_{0}\}}{1-\pi},$
where $m_{a}(\boldsymbol{W})=\operatorname{E}(Y|\boldsymbol{W},A=a)=\operatorname{E}\{Y(a)|\boldsymbol{W}\}$, $a=0,1$. Although $b_{\text{opt}}$ is unknown, it can be estimated from trial data using a regression model for $\operatorname{E}(Y|A,\boldsymbol{W})$ or a machine learning algorithm. Let $\widehat{b}$ be a generic estimator of $b_{\text{opt}}$ and assume that $\widehat{b}$ converges to some limit function $b$, which may be different from $b_{\text{opt}}$. Under quite general conditions, $\widehat{\delta}_{\text{aug}}(\widehat{b})$ (or a cross-fitted version of it) is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ with the same asymptotic variance $\sigma^{2}(b)$. This result holds regardless of the convergence rate of $\widehat{b}$, which supports the use of nonparametric machine learning methods to estimate $b_{\text{opt}}$.
Remark 1.
Covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function $b_{\text{opt}}$. Not only do we know that $b_{\text{opt}}$ minimizes the asymptotic variance $\sigma^{2}(b)$ of an augmented estimator, we can show that the value of $\sigma^{2}(b)$ is directly related to the mean square (i.e., $L_{2}$) distance between $b$ and $b_{\text{opt}}$. Write $\|d\|_{2}^{2}=\operatorname{E}\{d(\boldsymbol{W})^{2}\}$ for any real-valued function $d$ of $\boldsymbol{W}$. In Appendix, we show that, for any augmentation function $b$,
$\sigma^{2}(b)=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\|b-b_{\text{opt}}\|_{2}^{2}.$ (3)
In other words, the extra variance of a sub-optimal estimator, $\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})$, is proportional to the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$. The practical implication is that, while it may be unrealistic to hope that $b=b_{\text{opt}}$, it is generally helpful to reduce the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$.
3 Stratified Randomization
3.1 Description and Known Results
Stratified randomization involves stratifying the patient population on the basis of one or more baseline covariates, which we assume are included in $\boldsymbol{W}$. Let the support of $\boldsymbol{W}$ be partitioned as $\mathcal{W}=\cup_{k=1}^{K}\mathcal{W}_{k}$ with $\mathcal{W}_{k_{1}}\cap\mathcal{W}_{k_{2}}=\emptyset$ for $k_{1}\not=k_{2}$. Let $S=s(\boldsymbol{W})=\sum_{k=1}^{K}kI(\boldsymbol{W}\in\mathcal{W}_{k})$, and let $S_{i}=s(\boldsymbol{X}_{i})$, $i=1,\dots,n$. While simple randomization would be sufficient to ensure that $\operatorname{P}(A=1|S)=\operatorname{P}(A=1)=\pi$, stratified randomization aims to achieve better balance within each stratum defined by $S$. It does so by randomizing subjects in consecutive blocks within each stratum, with the constraint that the proportion of subjects with $A_{i}=1$ must be equal to $\pi$ within each completed block. Under stratified randomization, it is evident that $(A_{1},\dots,A_{n})$ are conditionally independent of $(\boldsymbol{Z}_{1},\dots,\boldsymbol{Z}_{n})$ given $(S_{1},\dots,S_{n})$. The observed data, $\{\boldsymbol{O}_{i}=(\boldsymbol{W}_{i},A_{i},Y_{i}),i=1,\dots,n\}$, are no longer independent across subjects because the $A_{i}$’s may depend on each other. On the other hand, the $\boldsymbol{O}_{i}$’s are identically distributed and their (common) marginal distribution is the same as the distribution of $\boldsymbol{O}=(\boldsymbol{W},A,Y)$ (described at the beginning of Section 2.2).
It is straightforward to deduce the asymptotic properties of augmented estimators from the recent theoretical work on CAR 17, 21, 22. As before, suppose $\widehat{b}$ is an estimator of $b_{\text{opt}}$ that converges to some fixed function $b$. Under stratified randomization, the augmented estimators $\widehat{\delta}_{\text{aug}}(b)$ and $\widehat{\delta}_{\text{aug}}(\widehat{b})$ are consistent and asymptotically normal with the same asymptotic variance, which we denote by $\sigma_{\text{st}}^{2}(b)$; that is, both $\sqrt{n}\{\widehat{\delta}_{\text{aug}}(b)-\delta\}$ and $\sqrt{n}\{\widehat{\delta}_{\text{aug}}(\widehat{b})-\delta\}$ converge to the same normal distribution with mean 0 and variance $\sigma_{\text{st}}^{2}(b)$. Wang et al. 21 give the following characterization of $\sigma_{\text{st}}^{2}(b)$:
$\sigma_{\text{st}}^{2}(b)=\sigma^{2}(b)-\frac{1}{\pi(1-\pi)}\operatorname{E}\left\{\left(\operatorname{E}[(A-\pi)\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}|S]\right)^{2}\right\}.$ (4)
Recall that $\sigma^{2}(b)$ is the common asymptotic variance of $\widehat{\delta}(b)$ and $\widehat{\delta}(\widehat{b})$ under simple randomization. Clearly, $\sigma_{\text{st}}^{2}(b)\leq\sigma^{2}(b)$, where equality holds if and only if $\operatorname{E}[(A-\pi)\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}|S]\equiv 0$. Thus, compared to simple randomization, stratified randomization does not increase, and typically decreases, the asymptotic variance of an augmented estimator.
3.2 Connection with Covariate Adjustment
In this subsection, we give a geometric interpretation to equation (4) that helps make a connection with covariate adjustment. We will show that the variance reduction effect of stratified randomization is equivalent to moving a given augmentation function $b$ closer to $b_{\text{opt}}$ under simple randomization. Recall from Remark 1 that, under simple randomization, a shorter distance to $b_{\text{opt}}$ leads to a smaller asymptotic variance.
Some notations are necessary for describing the result. We write $\mathcal{B}$ for the space of augmentation functions, i.e., =$\mathcal{B}=\{b:\|b\|_{2}<\infty\}$. This will be regarded as a Hilbert space with inner product $\langle b_{1},b_{2}\rangle=\operatorname{E}\{b_{1}(\boldsymbol{W})b_{2}(\boldsymbol{W})\}$. Let $\mathcal{C}$ denote the linear subspace of $\mathcal{B}$ consisting of functions of $S$; this can be formally defined as $\mathcal{C}=\{c\in\mathcal{B}:\operatorname{Var}\{c(\boldsymbol{W})|S\}\equiv 0\}$. Let $\mathcal{C}^{\perp}$ denote the orthogonal complement of $\mathcal{C}$ in $\mathcal{B}$, so $\mathcal{C}^{\perp}=\{b\in\mathcal{B}:\operatorname{E}\{b(\boldsymbol{W})|S\}\equiv 0\}$. These subspaces are illustrated in Figure 1. Let the operator $\Lambda$ be defined by $\Lambda b=b_{\text{opt}}+\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})$, where $\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})$ is the projection of $b-b_{\text{opt}}$ into $\mathcal{C}^{\perp}$. As shown in Figure 1, $\Lambda b$ is the result of projecting $b$ into the affine subspace $b_{\text{opt}}+\mathcal{C}^{\perp}$.
Figure 1: A geometric interpretation of stratified randomization (see Section 3.2 for definitions).
In Appendix, we show that, for any $b\in\mathcal{B}$,
$\sigma_{\text{st}}^{2}(b)=\sigma^{2}(\Lambda b)=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\left\|\Pi\left(b-b_{\text{opt}}|\mathcal{C}^{\perp}\right)\right\|_{2}^{2}.$ (5)
Thus, for an estimator with (asymptotic) augmentation function $b$, stratified randomization amounts to replacing $b$ with $\Lambda b$ under simple randomization. Because $\Lambda$ projects into a subspace containing $b_{\text{opt}}$, $\Lambda b$ is generally closer to $b_{\text{opt}}$ than $b$ is. By Remark 1, this implies $\sigma^{2}(\Lambda b)\leq\sigma^{2}(b)$.
Remark 2.
Equation (5) makes it clear that $\sigma_{\text{st}}^{2}(b)$ reaches its minimum, $\sigma^{2}(b_{\text{opt}})$, when and only when $b-b_{\text{opt}}\in\mathcal{C}$. In other words, $\sigma_{\text{st}}^{2}(b)$ is minimized by taking $b(\boldsymbol{W})=b_{\text{opt}}(\boldsymbol{W})+c(S)$ for an arbitrary function $c$. This result indicates that stratified randomization provides some latitude in efficient estimation by allowing $b$ to differ from $b_{\text{opt}}$ by a function of $S$. The fact that $\sigma_{\text{st}}^{2}(b)$ and $\sigma^{2}(b)$ have the same minimum value is consistent with the main result of Rafi 30, which states that, for estimating the average treatment effect, the same efficiency bound holds under CAR and simple randomization.
Remark 3.
Suppose we take $\widehat{\delta}_{\text{aug}}(b)$ as an initial estimator and seek to improve its efficiency under simple randomization by incorporating information from $S$. Under the augmentation approach, this can be done by subtracting from $\widehat{\delta}_{\text{aug}}(b)$ an augmentation term based on $S$, leading to
$\widehat{\delta}_{\text{aug}}(b)-\frac{1}{n}\sum_{i=1}^{n}(A_{i}-\pi)c(S_{i})=\widehat{\delta}_{\text{emp}}-\frac{1}{n}\sum_{i=1}^{n}(A_{i}-\pi)\{b(\boldsymbol{W}_{i})+c(S_{i})\}=\widehat{\delta}_{\text{aug}}(b+c),$
with asymptotic variance $\sigma^{2}(b+c)$ under simple randomization. According to Remark 1,
$\sigma^{2}(b+c)=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\|b+c-b_{\text{opt}}\|_{2}^{2},$
and its minimum value over $c$ is exactly equal to $\sigma_{\text{st}}^{2}(b)$:
$\min_{c\in\mathcal{C}}\sigma^{2}(b+c)$ $=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\min_{c\in\mathcal{C}}\|b+c-b_{\text{opt}}\|_{2}^{2}$ (6)
$=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\left\|\Pi\left(b-b_{\text{opt}}|\mathcal{C}^{\perp}\right)\right\|_{2}^{2}=\sigma_{\text{st}}^{2}(b),$
where the second step follows from geometry and the last one from equation (5). Thus, for a generic augmented estimator, stratified randomization is asymptotically equivalent to attaching an optimal augmentation term based on $S$ under simple randomization.
As an example, consider the empirical estimator $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$. With $b=0$, Remark 3 implies that $\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(c)$. Minimizing $\sigma^{2}(c)$ over $c\in\mathcal{C}$ is a special case of the efficient estimation problem considered in Section 2.2, with $\boldsymbol{W}$ replaced by $S$. Because $S$ is discrete, the minimum of $\sigma^{2}(c)$ is attained by the regression estimator based on model (1) with $\boldsymbol{V}=(I(S=2),\dots,I(S=K))^{\text{T}}$. With this choice of $\boldsymbol{V}$, model (1) is saturated and trivially correct, leading to a regression estimator that is efficient for $S$, which we denote by $\widehat{\delta}_{\text{reg}}^{S}$. Now, Remark 3 implies that the asymptotic variance of $\widehat{\delta}_{\text{emp}}$ under stratified randomization is equal to the asymptotic variance of $\widehat{\delta}_{\text{reg}}^{S}$ under simple randomization. This asymptotic equivalence makes intuitive sense. Stratified randomization produces balanced strata by design, while covariate adjustment can be regarded as an attempt to balance covariates by analysis. When covariate adjustment is focused on $S$ and optimally conducted, as in $\widehat{\delta}_{\text{reg}}^{S}$, it is expected to have the same asymptotic effect as stratified randomization based on $S$. It should be noted that this equivalence between covariate adjustment and stratified randomization is limited to the use of $S$. Covariate adjustment is able to incorporate information from additional baseline covariates of all types, whereas stratified randomization is constrained by the (typically small) number of strata that can be accommodated.
3.3 Practical Implications
In this subsection, we examine the implications of the results in Section 3.2 on clinical trial analysis and design. It is well recognized that stratified randomization should be acknowledged in the statistical analysis of trial data, which is usually interpreted to mean that treatment effect estimation should adjust for the stratification factor as a covariate 24, 25, 26. The recent FDA guidance on covariate adjustment 1 recommends that, under stratified randomization, “A covariate adjustment model should generally include strata variables and can also include covariates not used for stratifying randomization.” Accordingly, we will evaluate and compare the potential benefits of adjusting for the stratification factor and possibly other baseline covariates.
Let us start with the common practice of adjusting only for the stratification factor $S$. There are different ways to adjust for $S$, some of which may be more efficient than others. We will focus on the regression estimator $\widehat{\delta}_{\text{reg}}^{S}$, which makes optimal adjustment for $S$, as noted in Section 3.2. This implies that $\widehat{\delta}_{\text{reg}}^{S}$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(c_{\text{opt}})$, where $c_{\text{opt}}=\operatorname{arg\;min}_{c\in\mathcal{C}}\sigma^{2}(c)$ is the optimal augmentation function based on $S$ only. Under simple randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is generally more efficient than $\widehat{\delta}_{\text{emp}}=\widehat{\delta}_{\text{aug}}(0)$ because $\sigma^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c)\leq\sigma^{2}(0)$. Under stratified randomization, equation (6) implies that $\widehat{\delta}_{\text{emp}}$ has asymptotic variance $\sigma_{\text{st}}^{2}(0)=\min_{c\in\mathcal{C}}\sigma^{2}(0+c)=\sigma^{2}(c_{\text{opt}})$ and that $\widehat{\delta}_{\text{reg}}^{S}$ has asymptotic variance $\sigma_{\text{st}}^{2}(c_{\text{opt}})=\min_{c\in\mathcal{C}}\sigma^{2}(c_{\text{opt}}+c)=\sigma^{2}(c_{\text{opt}})$. Thus, under stratified randomization, $\widehat{\delta}_{\text{reg}}^{S}$ is as (in)efficient as $\widehat{\delta}_{\text{emp}}$ is. This is not surprising because the information in $S$ has already been used (optimally) in the randomization procedure. Only when we incorporate additional information from $\boldsymbol{W}$ into the analysis can we hope to improve efficiency over $\widehat{\delta}_{\text{emp}}$ under stratified randomization.
The information in $\boldsymbol{W}$ can be incorporated using a regression or augmented estimator described in Section 2.2. When all of $\boldsymbol{W}$ is included in covariate adjustment, the minimum attainable asymptotic variance is, according to Remark 2, $\sigma^{2}(b_{\text{opt}})=\min_{b\in\mathcal{B}}\sigma^{2}(b)$. Because $\mathcal{B}\supset\mathcal{C}$, this minimum is generally smaller than $\min_{c\in\mathcal{C}}\sigma^{2}(c)=\sigma^{2}(c_{\text{opt}})$, the minimum asymptotic variance that could be attained by adjusting for $S$ alone. The two minima are equal only when $\boldsymbol{W}$ contains no new prognostic information beyond $S$, which is unlikely given how $\boldsymbol{W}$ is chosen in the first place. Of course, $\sigma^{2}(b_{\text{opt}})$ is only a minimum and not necessarily the actual asymptotic variance of the estimator in use. However, in our numerical experience, it is difficult to find a situation in which an estimator that adjusts for all of $\boldsymbol{W}$ underperforms an estimator based on $S$, say $\widehat{\delta}_{\text{reg}}^{S}$.
In designing a trial with stratified randomization, one often faces difficult choices in terms of which covariates to include in the stratification and how many strata to use. Including more covariates increases the potential benefit of stratified randomization, but may result in a large number of strata, some of which may be small and difficult to balance. Some biomarkers may be time-consuming to ascertain, and their inclusion in stratified randomization may cause a delay in randomization and treatment. Fortunately, Remark 3 indicates that the efficiency benefit of stratified randomization can be recovered through covariate adjustment. Therefore, it is desirable but not essential to include all important covariates in the stratification. To illustrate this point, we give some examples of covariate-adjusted estimators that have the same asymptotic variance under simple versus stratified randomization. The regression estimator $\widehat{\delta}_{\text{reg}}$ has this property if $\pi=1/2$ or if the vector $\boldsymbol{V}$ in model (1) includes the dummy variables $(I(S=2),\dots,I(S=K))$, according to Corollary 1 of Wang et al. 21. (There can, and should, be other covariates in $\boldsymbol{V}$ to improve efficiency.) A generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ can be modified to have this property using the following calibration technique proposed by Bannick et al. 22. For $k=1,\dots,K$, define
$\widehat{c}_{k}=\frac{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)\widehat{\psi}_{(}\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=1,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)\widehat{\psi}(\boldsymbol{O}_{i})}{\sum_{i=1}^{n}I(A_{i}=0,S_{i}=k)}-\frac{\sum_{i=1}^{n}I(S_{i}=k)\widehat{b}(\boldsymbol{W}_{i})}{\sum_{i=1}^{n}I(S_{i}=k)},$
where
$\widehat{\psi}(\boldsymbol{O})=g^{\prime}(\overline{Y}_{1})\frac{A(Y-\overline{Y}_{1})}{\pi}-g^{\prime}(\overline{Y}_{0})\frac{(1-A)(Y-\overline{Y}_{0})}{1-\pi}.$
Let $\widehat{c}(S)=\sum_{k=1}^{K}\widehat{c}_{k}I(S=k)$; then $\widehat{\delta}_{\text{aug}}(\widehat{b}+\widehat{c})$ is the calibrated version of $\widehat{\delta}_{\text{aug}}(\widehat{b})$ with the desired property. (We note that $\widehat{c}$ estimates the function $c$ that attains the minima in equation (6), where $b$ is the limit of $\widehat{b}$.)
Remark 4.
Under stratified randomization, the calibration technique does not improve efficiency (as can be argued along the lines of Remark 3) but can be useful for variance estimation. The variance reduction effect of stratified randomization implies that variance estimators developed for simple randomization may be biased upward under stratified randomization, unless the point estimator of $\delta$ has the same asymptotic variance under simple versus stratified randomization, as is the case for a calibrated estimator. For a generic augmented estimator $\widehat{\delta}_{\text{aug}}(\widehat{b})$ with $\widehat{b}\to b$, let $\widehat{\delta}_{\text{aug}}(\widehat{b}+\widehat{c})$ be defined as in the preceding paragraph; then $\sigma_{\text{st}}^{2}(b)$ is the common asymptotic variance for the two estimators under stratified randomization, and is consistently estimated by
$\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{\psi}(\boldsymbol{O}_{i})-(A_{i}-\pi)\{\widehat{b}(\boldsymbol{W}_{i})+\widehat{c}(S_{i})\}\right]^{2}.$
If $\widehat{b}$ is obtained using a machine learning method, the above variance estimator may suffer from a re-substitution bias in finite samples, which can be removed using cross-validation; see Section 2.3 of Zhang and Ma 10. For a non-augmented estimator, one may appeal to its asymptotic equivalence to an augmented estimator (say $\widehat{\delta}_{\text{aug}}(b)$), find a consistent estimator of $b$ (say $\widehat{b}$), and then apply the calibration technique to obtain a consistent variance estimator under stratified randomization.
4 Simulation
To complement the preceding discussion based on asymptotic theory, this section reports a simulation study that evaluates the impact of covariate adjustment and stratified randomization in finite samples. In this simulation study, $\boldsymbol{W}=(W_{1},W_{2},W_{3})^{\text{T}}$ follows the trivariate standard normal distribution with zero correlations, and $S=s(\boldsymbol{W})=1+I(W_{1}>0)+2I(W_{2}>0)$ defines four equal-sized strata based on the signs of $W_{1}$ and $W_{2}$. With $\pi=1/2$, treatment assignment is either completely random (as in simple randomization) or stratified by $S$ with a fixed block size of four. The outcome $Y$ may be continuous or binary, and its dependence on $(\boldsymbol{W},A)$ may be described as follows:
Scenario A: $g^{\dagger}\{\operatorname{E}(Y|\boldsymbol{W},A)\}=-1+W_{1}-W_{2}+W_{3}+A(1+W_{2}-2^{-1}W_{3})$;
Scenario B: $g^{\dagger}\{\operatorname{E}(Y|\boldsymbol{W},A)\}=-1+W_{1}-W_{2}+W_{1}W_{3}-W_{2}W_{3}+A(1+W_{1}W_{2})$;
Scenario C: $g^{\dagger}\{\operatorname{E}(Y|\boldsymbol{W},A)\}=-1+W_{1}+W_{1}^{2}-(W_{2}\vee 0)^{2}+A(2+W_{3}-W_{3}^{2})$;
Scenario D: $g^{\dagger}\{\operatorname{E}(Y|\boldsymbol{W},A)\}=-1+\sqrt{W_{1}^{2}+W_{2}^{2}}-|W_{3}|+A$,
where $\vee$ denotes maximum and $g^{\dagger}$ is either the identity function (for continuous $Y$) or the logit function (for binary $Y$). In the continuous case, $Y$ is generated as $Y=\operatorname{E}(Y|\boldsymbol{W},A)+\varepsilon$, where $\varepsilon\sim N(0,1)$ independently of $(\boldsymbol{W},A)$. Scenario A follows a typical generalized linear model structure with interactions between $A$ and $\boldsymbol{W}$. Scenarios B and C feature additional interactions and non-linearity, and Scenario D represents further departure from the typical structure. In each scenario, 2,500 trials are simulated for each type of outcome (continuous or binary) and each randomization scheme (simple or stratified). Each simulated trial consists of $n=200$ or 500 subjects.
Let us first consider the case where $Y$ is continuous. In this case, the estimand $\delta$ is the mean difference between $Y(1)$ and $Y(0)$, and its true value is equal to 1 in each of the four scenarios described earlier. We compare six estimators of $\delta$: the empirical estimator $\widehat{\delta}_{\text{emp}}$, three regression estimators with different choices of $\boldsymbol{V}$, two augmented estimator (with or without calibration) where $\widehat{b}$ is obtained using a super learner 10. In the regression estimators, $h$ is the identity function (corresponding to ANCOVA) and $\boldsymbol{V}$ may be a vector of dummy variables for $S$, the original covariate vector $\boldsymbol{W}$, or $\boldsymbol{W}$ concatenated with dummy variables for $S$. The super learner in the augmented estimators is a linear combination (with coefficients optimized in five-fold cross-validation) of the following four algorithms: linear regression, linear regression with pairwise interactions, additive model, and recursive partitioning and regression tree. Under simple randomization, variance estimates are obtained using estimated influence functions, with the use of another five-fold cross-validation for the augmented estimators 10. Under stratified randomization, no adjustment in variance estimation is required for the regression estimators 21, and calibrated variance estimators (see Remark 4) are used for the empirical and augmented estimators. The different methods and designs (i.e., simple vs stratified randomization) are compared in terms of empirical bias, standard deviation, median standard error, and empirical coverage of 95% confidence intervals. The empirical method combined with simple randomization serves as a reference in efficiency comparisons. The relative efficiency of any method-design combination is defined as the inverse ratio of its empirical variance to that of the reference combination.
Tables 1 and 2 report simulation results for continuous $Y$ with $n=200$ or 500, respectively. The results indicate that estimation bias is negligible (relative to variability), that sampling variability is estimated accurately, and that empirical coverage is adequate, for any method-design combination. Therefore, we will focus on efficiency in comparing different estimators and designs. Under simple randomization, the empirical estimator typically has the lowest efficiency, and the other estimators may or may not bring a substantial improvement, depending on the scenario. The regression estimator adjusting for $S$ only improves efficiency if S is associated with $Y$, as is clearly the case in Scenarios A and B but not in Scenario D. The regression estimators that adjust for $\boldsymbol{W}$ (with or without $S$) improve efficiency (by a larger amount than that achieved by adjusting for $S$ only) in Scenarios A–C but not in Scenario D (where the ANCOVA models are severely mis-specified). The augmented estimators perform similarly to the regression estimators that adjust for $\boldsymbol{W}$ in Scenario A (where the ANCOVA models are correct) and outperform all other estimators in Scenarios B–D. Calibration has no appreciable impact on the super learner based augmentation method. Similar patterns are observed under stratified randomization with one notable exception: adjusting for $S$ only is no longer beneficial, which should be expected because stratified randomization has already made use of stratum information. In terms of utilizing stratum information, stratified randomization seems slightly more effective than covariate adjustment (with the regression estimator based on $S$) in small samples (e.g., $n=200$), although the difference is small and vanishes with increasing $n$. For the regression and augmented estimators based on $\boldsymbol{W}$, it is theoretically possible that stratified randomization may reduce the asymptotic variance, but simulation results indicate that such an improvement tends to be small in magnitude. This is consistent with the observation that calibration brings little improvement to the uncalibrated augmented estimator (under either randomization scheme). Including all of $\boldsymbol{W}$ in covariate adjustment (through regression or, preferably, augmentation) appears to be the main driver of estimation efficiency. When all of $\boldsymbol{W}$ is included in covariate adjustment (through regression or augmentation), it seems difficult to achieve further improvement by making another use of the information in $S$—be it through regression, calibration, or stratified randomization.
In the case of binary $Y$, the estimand is taken to be the log odds ratio, and its true value is found numerically to vary between 0.66 and 0.86 in the four scenarios considered. We compare six estimation methods that are analogous to those compared in the continuous case with the following modifications: the link function has become the logit function, and the ANCOVA models have been replaced by logistic regression models. Except for those changes, the methods are implemented and compared in the same manner as before. Tables 3 and 4 report simulation results for binary $Y$ with $n=200$ or 500, respectively. The results in Tables 3 and 4 follow the same patterns as those in Tables 1 and 2, although in the binary case efficiency gains tend to be smaller in magnitude.
5 Example
We now illustrate our findings by analyzing real data from a National Institute of Neurological Disorders and Stroke (NINDS) trial of recombinant tissue plasminogen activator (rt-PA) for treating acute ischemic stroke 31. This trial enrolled 624 patients with acute ischemic stroke at eight sites, randomized them in a $1:1$ ratio to rt-PA or placebo within three hours of stroke onset, and evaluated clinical outcomes at 24 hours and three months post-treatment. This trial consisted of two (temporally ordered) parts, with 291 patients enrolled in part 1 and 333 in part 2. The main difference between the two parts was that the time from stroke onset to treatment (TOT) was within 90 minutes in part 1 and between 90 minutes and 180 minutes in part 2. Except for that difference, patients in the two parts were treated and evaluated in the same manner. Within each part, the randomization was stratified by site using permuted blocks of various sizes. In our retrospective analysis, we combine the two parts and consider the overall randomization scheme as stratified by site and study part.
Following previous analyses of the NINDS trial data 31, 32, we compare treatments with respect to clinical outcomes at three months, as measured by the Barthel Index, the Modified Rankin Scale, the Glasgow Outcome Scale, and the National Institutes of Health Stroke Scale (NIHSS). The Barthel Index measures the ability to perform activities of daily living and yields a score that ranges from 0 to 100 (complete independence). The Modified Rankin Scale is a simplified overall assessment of function, with a score that ranges from 0 (absence of symptoms) to 5 (severe disability). The Glasgow Outcome Scale is a global assessment of function, with a score that ranges from 1 (good recovery) to 5 (death). The NIHSS is a 42-point scale with 0 indicating no stroke symptoms and higher scores for more severe symptoms. For ease of interpretation and comparison, each original score is linearly re-scaled to the unit interval in such a way that higher scores represent desirable outcomes. For each instrument, the efficacy of rt-PA relative to placebo is measured by the between-group difference (rt-PA minus placebo) in the re-scaled mean score.
In our analysis, the covariate vector $\boldsymbol{W}$ consists of continuous TOT, baseline NIHSS, age, preexisting disability, and history of diabetes. (Study part is not included in $\boldsymbol{W}$ because it is determined by TOT, and clinical site is not included because some sites have few patients.) Six patients with missing covariate data are excluded from this analysis. The remaining data is analyzed using the empirical method, the regression method based on $\boldsymbol{W}$, and the augmentation method based on $\boldsymbol{W}$ with or without calibration. (It is infeasible to adjust for strata in the regression method due to data sparsity.) The methods are identical to those compared in the simulation study (for a continuous outcome), except that the super learner library includes random forest in addition to the algorithms described in Section 4.
Table 5 shows point estimates of re-scaled mean score differences together with standard errors that do or do not account for stratified randomization. Accounting for stratified randomization results in minimal changes in the standard errors, suggesting that the stratification factor is not strongly prognostic. Adjusting for the other covariates ($\boldsymbol{W}$) leads to larger reductions in the standard errors. Taken together, the results in Table 5 clearly indicate that the use of rt-PA has a beneficial effect, which is statistically significant at the one-sided 2.5% level (without adjusting for multiplicity) for all scores but NIHSS.
6 Discussion
With focus on statistical efficiency, this article highlights a connection between covariate adjustment and stratified randomization with important practical implications. Specifically, stratified randomization can be regarded as a special form of covariate adjustment that makes optimal use of stratum information. Covariate adjustment is more general and able to incorporate information from all baseline covariates associated with clinical outcomes, although its effectiveness depends on the appropriateness of the method. In the analysis of trial data following stratified randomization, adjusting for the stratification factor only provides no efficiency benefit, and the key to efficient estimation is incorporating new prognostic information from other covariates. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. These findings are supported by both theoretical and empirical results.
While we do not focus on comparing different methods for covariate adjustment, our simulation results suggest that machine learning methods for covariate adjustment perform similarly to or better than standard methods based on parametric models, depending on the appropriateness of the parametric models. In reality, parametric models are almost certainly mis-specified unless they are saturated, which can happen only when all covariates are discrete and their joint distribution has a small number of possible values (relative to the sample size).
While we do not consider other forms of CAR in this article, we note that our findings are equally applicable to the biased coin design. This is because equation (4), the basis for our findings concerning stratified randomization, also holds for the biased coin design, according to Wang et al. 21.
While we do not consider hypothesis testing in this article, we note that the large-sample performance of a Wald test is largely driven by the asymptotic properties of the point estimator and the variance estimator used in the test. For example, if the point estimator is consistent and asymptotically normal and the variance estimator is biased upward, the Wald test is expected to be conservative (with a deflated type I error rate) in large samples. Assuming that no bias exists in variance estimation, the power of the test generally increases with the efficiency of the point estimator.
Appendix: Proofs
Proof of Equation (3)
It can be shown as in Web Appendix C of Zhang, Tsiatis and Davidian 4 that $(A-\pi)b_{\text{opt}}(\boldsymbol{W})$ is the projection of $\psi(\boldsymbol{O})\in L_{2}(\boldsymbol{O})$ into the linear subspace $\{(A-\pi)d(\boldsymbol{W}):d\in L_{2}(\boldsymbol{W})\}$. Therefore, the residual $\psi(\boldsymbol{O})-(A-\pi)b_{\text{opt}}(\boldsymbol{W})\}$ is orthogonal to $(A-\pi)d(\boldsymbol{W})$ for any $d\in L_{2}(\boldsymbol{W})$. In particular, we have
$\operatorname{Cov}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{opt}}(\boldsymbol{W})\},(A-\pi)(b-b_{\text{opt}})(\boldsymbol{W})\}=0.$
It follows that
$\sigma^{2}(b)$ $=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}$
$=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{opt}}(\boldsymbol{W})-(A-\pi)(b-b_{\text{opt}})(\boldsymbol{W})\}$
$=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{opt}}(\boldsymbol{W})\}+\operatorname{Var}\{(A-\pi)(b-b_{\text{opt}})(\boldsymbol{W})\}$
$=\sigma^{2}(b_{\text{opt}})+\operatorname{E}[\{(A-\pi)(b-b_{\text{opt}})(\boldsymbol{W})\}^{2}]$
$=\sigma^{2}(b_{\text{opt}})+\operatorname{E}\{(A-\pi)^{2}\}\operatorname{E}\{(b-b_{\text{opt}})(\boldsymbol{W})^{2}\}$
$=\sigma^{2}(b_{\text{opt}})+\pi(1-\pi)\|b-b_{\text{opt}}\|_{2}^{2}.$
Proof of Equation (5)
Let us first reproduce equation (4):
$\sigma_{\text{st}}^{2}(b)=\sigma^{2}(b)-\frac{1}{\pi(1-\pi)}\operatorname{E}\left\{\left(\operatorname{E}[(A-\pi)\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}|S]\right)^{2}\right\}.$
We note that
$\quad\operatorname{E}[(A-\pi)\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}|\boldsymbol{W}]$
$=\operatorname{E}\{(A-\pi)\psi(\boldsymbol{O})|\boldsymbol{W}\}-\operatorname{E}\{(A-\pi)^{2}b(\boldsymbol{W})|\boldsymbol{W}\}$
$=\sum_{a=0}^{1}\operatorname{P}(A=a|\boldsymbol{W})(a-\pi)\operatorname{E}\{\psi(\boldsymbol{O})|A=a,\boldsymbol{W}\}-\operatorname{E}\{(A-\pi)^{2}\}b(\boldsymbol{W})$
$=\pi(1-\pi)\operatorname{E}\{\psi(\boldsymbol{O})|A=1,\boldsymbol{W}\}-\pi(1-\pi)\operatorname{E}\{\psi(\boldsymbol{O})|A=0,\boldsymbol{W}\}-\pi(1-\pi)b(\boldsymbol{W})$
$=\pi(1-\pi)\{b_{\text{opt}}(\boldsymbol{W})-b(\boldsymbol{W})\},$
where the last step follows from the definition of $b_{\text{opt}}$. Therefore,
$\operatorname{E}[(A-\pi)\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}|S]\\ =\pi(1-\pi)\operatorname{E}\{b_{\text{opt}}(\boldsymbol{W})-b(\boldsymbol{W})|S\}=\pi(1-\pi)\Pi(b_{\text{opt}}-b|\mathcal{C})(\boldsymbol{W})$
and
$\sigma_{\text{st}}^{2}(b)-\sigma^{2}(b)=-\pi(1-\pi)\|\Pi(b_{\text{opt}}-b|\mathcal{C})\|_{2}^{2}=-\pi(1-\pi)\|\Pi(b-b_{\text{opt}}|\mathcal{C})\|_{2}^{2}.$ (A.1)
It follows immediately from equation (3) that
$\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})=\pi(1-\pi)\|b-b_{\text{opt}}\|_{2}^{2}.$ (A.2)
Adding (A.1) and (A.2) yields
$\sigma_{\text{st}}^{2}(b)-\sigma^{2}(b_{\text{opt}})=\pi(1-\pi)\left(\|b-b_{\text{opt}}\|_{2}^{2}-\|\Pi(b-b_{\text{opt}}|\mathcal{C})\|_{2}^{2}\right)=\pi(1-\pi)\left\|\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})\right\|_{2}^{2}.$